OpenStudy (anonymous):

Solve for x. 3(2x - 3) + 4(2 - x) = 0

OpenStudy (anonymous):

What is the range of the function y = x 2?

OpenStudy (anonymous):

What are your answer choices.?

OpenStudy (anonymous):

You don't need answer choices for this, just answer it

OpenStudy (anonymous):

Distribute first: 6x - 9 + 8 - 4x = 0 Then combine like terms: 2x -1 = 0 Isolate x: 2x = 1 divide both sides by 2: x = 0.5

OpenStudy (anonymous):

Wait.? Is this YOUR question. ? no.? oh okay. t(-,-t)

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

\(\Large 3(2x - 3) + 4(2 - x) = 0 \Leftrightarrow 6x -9+8-4x=0 \\ \Large \Leftrightarrow 2x=1 \Rightarrow x=\frac{1}{2}\)

OpenStudy (anonymous):

x does not = 0.

OpenStudy (anonymous):

Jung you are overcomplicating things.

OpenStudy (anonymous):

We're just helping so take seat and get your tampon out of your butt nd put it in your vagina like it belongs. k thanks. :)

OpenStudy (anonymous):

You're not a genuis so I wouldnt be so cocky